Anh paid $12.09 for a pack of 6 pairs of sport socks. What was the price of a pair of sport socks?
Answer by math_tutor2020(3816)


x = price of one pair
6x = price of 6 pairs
6x = 12.09
x = 12.09/6
x = 2.015
x = 2.02 dollars

Each pair of socks costs about $2.02

Notice if x = 2.01 then 6x = 6*2.01 = 12.06
While x = 2.02 leads to 6x = 6*2.02 = 12.12
12.09 is between 12.06 and 12.12, so it helps confirm the answer.
We won't land on 12.09 exactly.